# Investigating the importance of social vulnerability in opioid-related mortality across the United States

### v2

@@ Acknowledgments
Funding
This work is sponsored by the US Department of Veterans Affairs using resources from the Knowledge Discovery Infrastructure which is located at the Oak Ridge National Laboratory and supported by the Office of Science of the U.S. Department of Energy. This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the US Department of Energy (DOE). The US government retains and the publisher, by accepting the article for publication, acknowledges that the US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for US government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan)
Author contributions
AD conducted all core research activities, curated and analyzed the data, generated the code, and wrote the original manuscript. AS and VM supervised the research while providing general and methodological expertise. DM suggested the use of Tobler for interpolating data in Connecticut. All authors reviewed the manuscript and contributed to revisions.

### v3

@@ Acknowledgments
CRediT authorship contribution statement
Andrew Deas: Writing – original draft, Writing – review and editing, Data curation, Methodology, Investigation, Formal analysis, Visualization, Software. Adam Spannaus: Writing – review and editing, Formal analysis, Methodology, Investigation, Supervision, Conceptualization, Funding acquisition. Dakotah Maguire: Writing – review and editing, Visualization, Methodology. Jodie Trafton: Writing – review and editing, Project administration, Funding acquisition. Anuj J. Kapadia: Writing – review and editing, Project administration, Funding acquisition, Supervision. Vasileios Maroulas: Writing – review and editing, Formal analysis, Methodology, Investigation, Supervision.
@@
Funding
This quality improvement initiative is sponsored by the U.S. Department of Veterans Affairs (VA), is supported by and utilizes VA-funded resources from the Knowledge Discovery Infrastructure (KDI) at Oak Ridge National Laboratory and is backed by the Department of Energy (DOE) Office of Science. The manuscript has been authored by UT-Battelle LLC under contract DE-AC05-00OR22725 with the DOE. The US government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce this manuscript or allow others to do so for US government purposes. DOE will provide public access to these results in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). The initiative aimed is to improve care for Veterans within the Veterans Health Administration (VHA) and involved models specifically tailored to the VHA system. Therefore, it is not generalizable outside of VHA and is officially considered non-research by the VHA. Oak Ridge National Laboratory received Institutional Review Board (IRB) approval due to the secondary use of patient data, which complies with internal policies and standards. The authors acknowledge the broader partnership and express gratitude to the Veterans receiving care at VA.